# Discriminative and Consistent Representation Distillation

### v4

@@ Abstract
Knowledge Distillation (KD) aims to transfer knowledge from a large teacher model to a smaller student model. While contrastive learning has shown promise in self-supervised learning by creating discriminative representations, its application in knowledge distillation remains limited and focuses primarily on discrimination, neglecting the structural relationships captured by the teacher model. To address this limitation, we propose Discriminative and Consistent Distillation (DCD), which employs a contrastive loss along with a consistency regularization to minimize the discrepancy between the distributions of teacher and student representations. Our method introduces learnable temperature and bias parameters that adapt during training to balance these complementary objectives, replacing the fixed hyperparameters commonly used in contrastive learning approaches. Through extensive experiments on CIFAR-100 and ImageNet ILSVRC-2012, we demonstrate that DCD achieves state-of-the-art performance, with the student model sometimes surpassing the teacher’s accuracy. Furthermore, we show that DCD’s learned representations exhibit superior cross-dataset generalization when transferred to Tiny ImageNet and STL-10111Code is available at https://github.com/giakoumoglou/distillers..
@@ 2.1 Knowledge Distillation
Logit-based Methods
@@
Feature-based Methods
@@
Architecture-aware Methods
@@ 3.2 Discriminative and Consistent Distillation
where $\phi(\mathbf{u},\mathbf{v})=\mathbf{u}^{T}\mathbf{v}/|\mathbf{u}||\mathbf{v}|$ represents the cosine similarity function, with $\tau$ serving as the temperature parameter, $b$ as a bias parameter, and $N$ as the total number of negatives. This formulation effectively transforms into a cross-entropy loss, where each student representation $\mathbf{z}_{i}^{S}$ must identify its corresponding teacher representation $\mathbf{z}_{i}^{T}$ among all other teacher representations in the batch. The objective essentially becomes a classification task: student embeddings must ”classify” their matching teacher embeddings correctly, with the normalized similarities acting as logits and positive pair indices as class labels. The parameters $\tau$ and $b$ provide fine-grained control over this classification process, determining its sharpness and scale respectively.
Consistent Distillation
@@ 3.3 Implementation Details
Memory-efficient Sampling
Instead of using a large memory buffer for contrasting representations as in CRD [66], we use the negative samples that naturally co-exist within the batch. This approach significantly reduces memory requirements while maintaining effective contrastive learning.
Adaptive Temperature Scaling
@@ 4.2 Main Results
4.3 Visualization
We present visualizations that analyze the learned representations and knowledge transfer across different distillation approaches.
Inter-class Correlations
@@ 4.3 Visualization
t-SNE Visualization
Figure 3 presents t-SNE [69] visualizations of embeddings from WRN-40-2 (teacher) and WRN-40-1 (student) on CIFAR-100. Compared to standard training and [74], DCD better aligns student embeddings with the teacher, preserving structural relationships in the feature space.
@@ 5 Conclusions
We have presented DCD, a knowledge distillation method that combines contrastive learning with consistency regularization to improve the traditional KD process. Our method achieves state-of-the-art performance through memory-efficient in-batch negative sampling and adaptive temperature scaling, eliminating the need for large memory banks while automatically tuning contrast levels during training. Through extensive experimentation across CIFAR-100, ImageNet, STL-10, and Tiny ImageNet datasets, we have demonstrated significant improvements over existing methods. Unlike previous methods such as CRD [66] that require large memory banks or WSLD [81] and IPWD [55] that focus solely on instance discrimination, our approach achieves superior performance while being more memory-efficient and capturing both local and global structural information. The effectiveness of DCD has been validated in both same-architecture and cross-architecture scenarios, with student models in several cases exceeding their teachers’ performance, particularly in WRN-40-2 to WRN-16-2 and WRN-40-2 to ShuffleNet-v1 configurations. We hope this work will inspire future research in knowledge distillation.
@@ Acknowledgments
We would like to express our gratitude to Andreas Floros for his valuable feedback, particularly his assistance with equations and notations. We also acknowledge the computational resources and support provided by the Imperial College Research Computing Service (http://doi.org/10.14469/hpc/2232), which enabled our experiments.

### v5

@@ Abstract
Knowledge Distillation (KD) aims to transfer knowledge from a large teacher model to a smaller student model. While contrastive learning has shown promise in self-supervised learning by creating discriminative representations, its application in knowledge distillation remains limited and focuses primarily on discrimination, neglecting the structural relationships captured by the teacher model. To address this limitation, we propose Discriminative and Consistent Distillation (DCD), which employs a contrastive loss along with a consistency regularization to minimize the discrepancy between the distributions of teacher and student representations. Our method introduces learnable temperature and bias parameters that adapt during training to balance these complementary objectives, replacing the fixed hyperparameters commonly used in contrastive learning approaches. Through extensive experiments on CIFAR-100 and ImageNet ILSVRC-2012, we demonstrate that DCD achieves state-of-the-art performance, with the student model sometimes surpassing the teacher’s accuracy. Furthermore, we show that DCD’s learned representations exhibit superior cross-dataset generalization when transferred to Tiny ImageNet and STL-10111Code is available at: https://github.com/giakoumoglou/distillers..
@@ 2.1 Knowledge Distillation
Logit-based methods.
@@
Feature-based methods.
@@
Architecture-aware methods.
@@ 3.2 Discriminative and Consistent Distillation
where $\phi(\mathbf{u},\mathbf{v})=\mathbf{u}^{\top}\mathbf{v}/|\mathbf{u}||\mathbf{v}|$ represents the cosine similarity function, with $\tau$ serving as the temperature parameter, $b$ as a bias parameter, and $N$ as the total number of negatives. This formulation effectively transforms into a cross-entropy loss, where each student representation $\mathbf{z}_{i}^{S}$ must identify its corresponding teacher representation $\mathbf{z}_{i}^{T}$ among all other teacher representations in the batch. The objective essentially becomes a classification task: student embeddings must ”classify” their matching teacher embeddings correctly, with the normalized similarities acting as logits and positive pair indices as class labels. The parameters $\tau$ and $b$ provide fine-grained control over this classification process, determining its sharpness and scale respectively.
Consistent distillation.
@@ 3.3 Implementation Details
Memory-efficient sampling.
Instead of using a large memory buffer for contrasting representations as in CRD [66], we use the negative samples that naturally co-exist within the batch. This approach significantly reduces memory requirements while maintaining effective contrastive learning. Our in-batch negative sampling eliminates the need to maintain and update large memory banks that store thousands of feature vectors, which can consume significant GPU memory It also simplifies the implementation by removing the complexity of memory bank management, including challenges related to feature staleness and queue maintenance [30]. Unlike memory banks that may contain stale features from earlier training iterations, our approach always uses the most recent representations within the current batch.
Learnable temperature.
@@ 4.3 Visualization
We present comprehensive visualizations that analyze the learned representations and knowledge transfer patterns across different distillation approaches, providing qualitative insights.
Inter-class correlations.
t-SNE visualization.
Figure 3 presents t-SNE [69] visualizations of embeddings from WRN-40-2 (teacher) and WRN-40-1 (student) on CIFAR-100. Compared to standard training and [74], DCD better aligns student embeddings with the teacher, preserving structural relationships in the feature space. The preservation of semantic relationships between class clusters demonstrates our method’s effectiveness in transferring the teacher’s knowledge organization.
@@ 5 Conclusions
We have presented DCD, a knowledge distillation method that combines contrastive learning with consistency regularization to improve the traditional KD process. Our method achieves state-of-the-art performance through memory-efficient in-batch negative sampling and adaptive temperature scaling, eliminating the need for large memory banks while automatically tuning contrast levels during training. Through extensive experimentation across CIFAR-100, ImageNet, COCO, STL-10, and Tiny ImageNet datasets, we have demonstrated significant improvements over existing methods. Unlike previous methods such as CRD [66] that require large memory banks or WSLD [81] and IPWD [55] that focus solely on instance discrimination, our approach achieves superior performance while being more memory-efficient and capturing both local and global structural information. The effectiveness of DCD has been validated in both same-architecture and cross-architecture scenarios, with student models in several cases exceeding their teachers’ performance. We hope this work will inspire future research in knowledge distillation.
@@ Acknowledgments
We would like to express our gratitude to Andreas Floros for his valuable feedback, particularly his assistance with equations, notations, and insightful discussions that greatly contributed to this work. We also acknowledge the computational resources and support provided by the Imperial College Research Computing Service (http://doi.org/10.14469/hpc/2232), which enabled our experiments.